OpenStudy (anonymous):

Please check my answers? I have a '*' next to my answer, can someone please please check them for me? 10. What is the algebraic expression for the following word phrase: the product of g and 4? a. g - 4 b. 4g* c. g/4 d. g + 4 What is the algebraic expression for the following word phrase: the quotient of 8 and the difference of x and m? a. 8/x-m b. 8 - x/m* c. x-m /8 d. 8-x / m What is the algebraic expression for the following word phrase: the quotient of x and 6y? a. x/6 b. x/6y c. 6xy* d. x - 6 and

OpenStudy (danjs):

for 2, group the numerators and denominaotors with parenthesis that have multiple terms

OpenStudy (danjs):

quotient means divide, quotient of A and B is A/B Product is multiply

OpenStudy (anonymous):

So, would number 2 actually be "A?"

OpenStudy (danjs):

quotient of 8 and the difference between x and m Division for the quotient, the difference of x and m is x-m the quotient of 8 and that difference is then 8 / (x-m)

OpenStudy (danjs):

yes, but as typed, you have to put parenthesis around (x-m) or it is not the same

OpenStudy (anonymous):

Alright, thank you! but what about the other two? Math is one of my worst subjects in school.

OpenStudy (danjs):

8/x - m = \[\frac{ 8 }{ x }-m\] group the denominators for one line typing, lik eyou would put it into a ti-83 or something

OpenStudy (danjs):

8/(x-m) =\[\frac{ 8 }{ x-m }\]

OpenStudy (danjs):

what is the third problem then?

OpenStudy (anonymous):

Alright so, the third problem would be "B" right? x/6y

OpenStudy (danjs):

yes, but remember to group the bottom , otherwise as you typed it would be x/6y = (x*y)/6